# VIPTR: A Vision Permutable Extractor for Fast and Efficient Scene Text Recognition

### v3

@@ 4.1 Further Analysis
To gain a deeper insight into VIPTR, we conducted controlled experiments on both IC13 (regular) and IC15 (irregular) under various configurations. The English dataset used in the experiment consists of COCO-Text and part of the real scene training set, with a total of 45054 images over 500 epochs of training. In addition, we conduct a series of comparative experiments to verify the effectiveness of our models on image texts of varying lengths. The Chinese Document Dataset is used for training and validation. Three test sets are employed: Testset1, the standard set from the Chinese Document Dataset; Testset2, consisting of 2,367 images cropped from real documents with lengths between 320 and 1500 pixels; and Testset3, featuring 638 images, also derived from real documents but with lengths exceeding 1500 pixels. For efficiency, all the experiments are carried out by using VIPTR-T without the rectification module and data augmentation.
Positional Embedding IC13 IC15 FLOPs (G)
None 76.7 59.7 0.25
APE 77.3 63.5 0.26
CPE 77.3 63.6 0.26
LePE 79.1 65 0.26
Table 4: Ablation study on Positional Embeddings.
@@
VIPTRv1-T 98.8 92.1 93.7
Table 5: Ablation studies on inputs of different lengths.
IC13 SVT IIIT5k IC15 SVTP CUTE
CRNN Shi et al. (2016) 91.1 81.6 82.9 69.4 70.0 65.5 78.5 53.4 8.3 6.3
Rosetta Borisyuk et al. (2018) 90.9 84.7 84.3 71.2 73.8 69.2 80.3 - 44.3 10.5
ViTSTR Atienza (2021) 93.2 87.7 88.4 78.5 81.8 81.3 85.6 - 85.5 11.2
SVTR-T Du et al. (2022) 96.3 91.6 94.4 84.1 85.4 88.2 90.7 67.9 6.0 4.5
SVTR-L Du et al. (2022) 97.2 91.7 96.3 86.6 88.4 95.1 92.8 72.1 40.8 18.0
ASTER Shi et al. (2018) 90.8 89.5 93.4 76.1 78.5 79.5 86.5 54.5 27.2 -
NRTR Sheng et al. (2019) 94.7 88.3 86.5 74.1 79.4 88.2 84.0 - 31.7 160.0
SAR Li et al. (2019) 91.0 84.5 91.5 69.2 76.4 83.5 83.6 62.5 57.5 120.0
SATRN Lee et al. (2020) 95.7 93.5 96.1 84.1 88.5 90.3 91.9 - - -
SRN Yu et al. (2020) 95.5 91.5 94.8 82.7 85.1 87.8 90.4 60.1 54.7 25.4
ABINet Fang et al. (2021) 97.4 93.5 96.2 86.0 89.3 89.2 92.6 - 36.7 51.3
VisionLAN Wang et al. (2021b) 95.7 91.7 95.8 83.7 86.0 88.5 91.2 - 32.8 28.0
VST Tang et al. (2021) 96.4 93.8 96.3 85.4 88.7 95.1 92.6 - 64.0 -
MATRN Na et al. (2022) 95.8 94.9 96.7 82.9 90.5 94.1 92.3 - 44.2 29.6
MaskOCR* Lyu et al. (2022) 97.7 93.7 95.4 86.6 89.0 87.5 92.5 68.6 36.0 24.1
MaskOCR Lyu et al. (2022) 96.8 94.7 95.3 87.1 89.3 90.6 92.7 68.8 100.0 -
TrOCR Li et al. (2023b) 98.3 93.2 91.0 84.0 91.0 89.6 90.2 - 558.0 20.8
AutoSTR Zhang et al. (2020) 94.2 90.9 94.7 81.8 81.7 84.0 89.5 - 10.4 207.0
PREN* Yan et al. (2021) 94.7 92.0 92.1 79.2 83.9 81.3 88.0 - 29.1 40.0
VIPTRv1-T (Tiny) 96.8 93.5 93.3 87.2 87.1 87.4 91.4 68.7 5.9 4.2
VIPTRv2-T 95.8 92.6 92.6 85.4 86.0 85.0 90.3 65.8 5.1 3.3
VIPTRv1-L (Large) 97.3 94.5 95.1 88.4 89.6 88.8 92.9 72.6 39.5 16.9
VIPTRv2-B (Base) 97.0 94.2 94.8 88.0 90.0 90.2 92.7 71.0 22.1 8.0
Table 6: Results on six English and Chinese scene benchmarks tested against existing methods. The model types from top to bottom are CTC-based, attention-based, LM-based, and others. * means the lightweight version of the corresponding model. The speed is the average inference time on one Tesla V100 GPU with batch size 4 (to align with Du et al. (2022)) for 3000 English image texts.
Effective Permutation of Attention Blocks
There are various ways to match the global and Local Attention blocks in each stage. In previous researchv Du et al. (2022), it was found that by placing the Local Attention block before the Global Attention block, it is beneficial to both feature mining of individual symbols and to guide the Global Attention block to focus on long-term dependency capture between symbol features, while effectively saving the inference cost of the model. The top half and the bottom half of Table 1 respectively show the performance of VIPTRv1-T and VIPTRv2-T composed of various Local and Global Attention and their permutations. $[L_{i}][L_{j}G_{s}][G_{t}]$ means for each stage, Local Attention blocks are carried out at first, and then Global Attention blocks. $[L_{i}][L_{j}//G_{s}][G_{t}]$ means in the stage 2-3, Local and Global Attention blocks are carried in parallel. The $G_{1}$ means MHSA, the $G_{2}$ means ORSA, the $L_{1}$ means CSWin and the $L_{2}$ means MaSA. Others are defined similarly.
It can be observed that almost every scheme gives a certain accuracy improvement compared to CRNN. We believe that the improvements are attributed to the comprehensive perception of character components features by the self-attention mechanism. The relatively large gains on irregular text further explain the self-attention mixing block is helpful to feature learning in complex scenarios. It is observed that $[L_{2}][L_{2}G_{2}][G_{1}]$ reports the best accuracy in the top half of Table 1 and $[L_{2}][L_{2}//G_{2}][G_{1}]$ achieves best in the bottom half of Table 1. The best one gives accuracy gains of 1.4% and 2.8% on IC13 and IC15 when compared with SVTR-T. It is observed that $[L_{1}][L_{1}//G_{2}][G_{1}]$ reports the fastest inferencing speed.
4.1.1 Effectiveness of Positional Embeddings.
As indicated in Table 4, various PE schemes exhibit slight variations in recognition accuracy. The LePE scheme surpasses the default one (None) by 2.4% and 5.3% on the two datasets, demonstrating its effectiveness, particularly in the case of irregular text.
4.1.2 Results for different lengths inputs testing
In Table 5, the performance of the three models on the three test sets is quite different. SVTR’s inference performance starts to collapse as the input length gets larger. VIPTR not only supports arbitrary lengths of image text input, but the inference performance will not decrease as the input length increases.
5 Conclusion
@@
In this study, we propose VIPTR, an innovative vision permutable extractor for fast and efficient Scene Text Recognition (STR). VIPTR excels at extracting character features and capturing contextual dependencies between character components, all while being unrestricted by input size during inference. These grants VIPTR notable accuracy, efficiency, and cross-lingual adaptability. We have also developed different versions of VIPTR to cater to a range of application requirements. Experiments on both English and Chinese benchmarks generally validate the effectiveness of our models. It achieves highly competitive or superior accuracy compared to state-of-the-art methods while maintaining faster processing speeds. Our future goal is to integrate VIPTR with an efficient text decoder to further elevate STR performance.

### v5

@@ 4.2. Ablation Study
To gain a deeper insight into SVIPTR, we conducted controlled experiments on both IC13 (regular) and IC15 (irregular) under various configurations. The training set and experimental Settings in the English scenario are unchanged. In addition, we conduct a series of comparative experiments to verify the effectiveness of our models on image texts of varying lengths. The Chinese Document Dataset is used for training and validation. Three test sets are employed: Testset1, the standard set from the Chinese Document Dataset; Testset2, consisting of 2,367 images cropped from real documents with lengths between 320 and 640 pixels; and Testset3, featuring 638 images, also derived from real documents but with lengths exceeding 640 pixels. For efficiency, all the experiments are carried out by using SVIPTR-T without the rectification module and data augmentation.
Effective Permutation of Attention Blocks. The top half and the bottom half of Table 1 respectively show the performance of SVIPTRv1-T and SVIPTRv2-T composed of various Local and Global Attention and their permutations. It can be observed that almost every scheme gives a certain accuracy improvement compared to CRNN. We believe that the improvements are attributed to the comprehensive perception of character components features by the self-attention mechanism. The relatively large gains on irregular text further explain the self-attention mixing block is helpful to feature learning in complex scenarios. It is observed that $[L_{2}][L_{2}G_{2}][G_{1}]$ reports the best accuracy in the top half of Table 1 and $[L_{2}][L_{2}//G_{2}][G_{1}]$ achieves best in the bottom half of Table 1. The best one gives accuracy gains of 1.4% and 2.8% on IC13 and IC15 when compared with SVTR-T. It is observed that $[L_{1}][L_{1}//G_{2}][G_{1}]$ reports the fastest inferencing speed. In general, SVIPTR has better performance than SVTR without using rectification module, and it is the fastest and highest performance visual feature encoder.
Effectiveness of Positional Embeddings. As indicated in Table 5, various PE schemes exhibit slight variations in recognition accuracy. The LePE scheme surpasses the default one (None) by 2.4% and 5.3% on the two datasets, demonstrating its effectiveness, particularly in the case of irregular text. The use of LePE not only improves the recognition accuracy without affecting the inference efficiency but also is insensitive to the input length, which enhances the versatility of SVIPTR.
Table 5. Ablation study on Positional Embeddings.
Model + Positional Embedding IC13 IC15 FLOPs (G)
SVIPTRv1-T + None 93.7 79.7 0.25
SVIPTRv1-T + APE 94 78.2 0.26
SVIPTRv1-T + CPE 94.3 78.9 0.26
SVIPTRv1-T + LePE 94.6 80.3 0.26
Results for different lengths inputs testing. In Table 6, the performance of the three models on the three test sets is quite different. SVTR’s inference performance starts to collapse as the input length gets larger. SVTR uses APE and SWin self-attention with static Windows, which does not have CNN-like translation invariance. So the length of inference images are limited by the size of training images. If inference images’ lengths are beyond the pre-set, resizing the image is required, which results in distortion. SVIPTR is not only length-insensitive, but the inference performance will not decrease as the input length increases.
Table 6. Ablation studies on inputs of different lengths.
@@
SVIPTRv1-T 98.8 92.1 93.7
4.3. Attention Visualization
We have visualized the attention maps of SVIPTR when decoding different character components. Finally, the proposed SVIPTRv2-T model without TPS and STN (It is obtained by training a large number of image texts with a uniform shape size of 32x96) with the weakest performance is selected to generate the attention map and shown in this paper, which further proves the high performance of all variants of SVIPTR. As shown in Figure 4, the picture in the first row is the original image (the shape size is $32\times 480$), and the right side is the model prediction result. The following attention map can be interpreted as the global and local attention mechanisms that play a role in the whole recognition process. The first attention map represents the degree of attention given to the character ”e” in the recognition process. From here, it can be seen that the model can maintain strong attention to the target character in a relatively long time span, which verifies the effectiveness of the global attention mechanism. The subsequent attention map is composed of selecting an appropriate timestamp corresponding to the attention of each character component in the original image. From the phenomenon that almost each character is covered by a bright color, it can be seen that the model allocates enough attention to the local vicinity of each character component, which verifies the effectiveness of the local attention mechanism. Moreover, the attention does not significantly decay as the character sequence becomes longer, which means that the dependencies between different character components are successfully captured, which further indicates that the model has good long-distance reasoning ability.
Figure 4 indicates that SVIPTRv2-T captures multiple features and cues locally, throughout, and across characters, which is consistent with the claim that SVIPTR senses multi-granularity character component features. It also verifies that the inference of SVIPTR is not sensitive to the length of the image, and the weakest performing variant also has strong long-distance inference ability. This again illustrates the effectiveness of SVIPTR and its variants.
Figure 4. Visualization of the attention maps for SVIPTRv2-T. The character in red means it is missed or an error.
[…]
Visualization of the attention maps for SVIPTRv2-T. The character in red means it is missed or an error.
5. Conclusion
@@
In this study, we propose SVIPTR, an innovative vision permutable extractor for fast and efficient Scene Text Recognition (STR). SVIPTR excels at extracting and fusing visual and semantic features of image text, and it is insensitive to the length of the inference input. These grants SVIPTR notable accuracy, efficiency, and cross-lingual adaptability. We have also developed different versions of SVIPTR to cater to a range of application requirements. Experiments on both English and Chinese benchmarks generally validate the effectiveness of our models. It achieves highly competitive or superior accuracy compared to state-of-the-art methods while maintaining faster processing speeds. Our future goal is to integrate SVIPTR with an efficient text decoder to further elevate STR performance.
Acknowledgements.
This work was supported in part by the National Natural Science Foundation of China (Grant Nos. U1636211, U2333205, 61672081, 62302025, 62276017), a fund project: State Grid Co., Ltd. Technology R&D Project (ProjectName: Research on Key Technologies of Data Scenario-based Security Governance and Emergency Blocking in Power Monitoring System, Proiect No.: 5108-202303439A-3-2-ZN), the 2022 CCF-NSFOCUS Kun-Peng Scientific Research Fund and the Opening Project of Shanghai Trusted Industrial Control Platform and the State Key Laboratory of Complex & Critical Software Environment (Grant No. SKLSDE-2021ZX-18).